OpenStudy (anonymous):

midpoints fllowing segment by pairs of endpoints (1,4) (6,-2)

ganeshie8 (ganeshie8):

use the midpoint formula :- midpoint connecting (x1, y1) (x2, y2) is, (x1+x2 /2 , y1+y2 /2)

ganeshie8 (ganeshie8):

(1, 4) (6, -2) x1 y1 x2 y2

ganeshie8 (ganeshie8):

plug them in the formula

OpenStudy (jdoe0001):

\(\bf \text{middle point of 2 points }\\ \left(\cfrac{x_2 + x_1}{2},\cfrac{y_2 + y_1}{2} \right)\\ (x_1, y_1) = (1,4)\qquad \qquad (x_2,y_2) = (6, -2)\)